Show me how to break this down please.
29% of adults would pay more for environmentally friendly products. You randomly select 10 adults. Find the probability that the number of adults who would pay more for environmentally friendly products is​ (a) exactly​ two, (b) more than​ two, and​ (c) between two and​ five, inclusive.

Answer by Boreal(15235)

exactly two means 10C2, the number of ways to choose 2, *.29^2*.71^8=0.2444
more than 2 is 1-0,1,2
probability of 0 is .71^10=0.0326
probability of 1 is 10*.29*.71^9=0.1330
probability of the 3 is 0.4100
probability of everything else, the answer, is 0.5900
between 2 and 5:
probability of 2 is 0.2444
of 3 is 10C3*0,29^3*0.71^7=0.2662
of 4 is 0.1903
of 5 is 0.0933
total is 0.7942